36. When the price of good X decreased by 10%, the quantity demanded for good Y increased by 15%. Which of the following statements is correct?

Goods X and Y are substitutes.

Good X is a normal good, and good Y is an inferior good.

Goods X and Y are complements.

Good X is an inferior good, and good Y is a normal good.

37. When the price of good X decreased by 10%, the quantity demanded for good Y increased by 15%. What is the cross-price elasticity of demand between goods X and Y?

–1.50

–0.67

0.67

1.50

38. When consumer income increased by 15%, the quantity demanded for a good increased by 25%. What is the income elasticity of demand?

–1.67

–0.60

0.60

1.67

39. Which of the following is NOT a characteristic of indifference curves?

They are infinite in number.

They never touch each other.

They never move.

Marginal utility is constant along an indifference curve.

They always slope down.

40. Which of the following is true about a budget line?

A budget line rotates when there is a change in the consumer’s budget.

A budget line shifts when there is change in the price of one of the goods/services the consumer purchases.

A budget line shows all of the combinations of two goods/services a consumer can afford when all of the budget is spent.

A budget line shows which combinations of goods/services a consumer finds equally desirable.

A budget line is graphed with a positive slope.
Answer #1

36) option A)

as price of X falls, Quantity Demanded of Y rises,

So two are Substitutes

37) option A)

Cross price Elasticity = +15/-10 = -1.5

38) option D)

Income Elasticity = %∆ in Q/%∆ in income

= 25/15

= 5/3 = 1.67

39) option. D)

MU is not Constant along IC,

Utility is Constant along IC

​​​​​

​​40) option C)

budget line Slope downwards , negative Slope

It shifts, when income changes,

It rotates when Price changes